OpenStudy (anonymous):

a 7-kg bowling ball moving at 8m/s strikes a 1-kg bowling pin and slows to 7m/s in .040 s. A)What is the deceleration of the bowling ball? How much force acts on thee bowling pin?

OpenStudy (loser66):

you have \(v_0 = 8 m/s\), \(\huge v_{final}= 7m/s\) time t = 0.040s . so?? just apply formula \(\huge a= \dfrac{ V_{final} - V_0}{t}\) to get the acceleration, the minus sign shows it is deceleration. that's it

OpenStudy (loser66):

when getting a, you have formula F = ma to get force, the unit is N (Newton)

OpenStudy (anonymous):

acceleration would be (7-8)/,040?

OpenStudy (anonymous):

a=-193

OpenStudy (anonymous):

@Loser66

OpenStudy (loser66):

7-8 = -1 and then divided by 0.040 I got -25m/s^2

OpenStudy (anonymous):

o yeah, my mistake

OpenStudy (anonymous):

so my answer for a would be how much it decelerated?

OpenStudy (anonymous):

Then I plug a into F=ma to find the forc acted on the bowling pin?

OpenStudy (anonymous):

@Loser66

OpenStudy (loser66):

1/ your answer should be : the deceleration is a = - 25 m/s^2 2/ yes, just plug that a into F = ma to calculate force F Q: is it not physics? why is it here? it's math field

OpenStudy (anonymous):

yes it is physics, but I usually get help faster in math.. and this is the math part of physics, right?(: Thanks anyways

OpenStudy (loser66):

hehehe, you are impish. ok, then